OpenStudy (anonymous):

I have an infinite serie sum from 1 to infinite of 1/( n*(n+1) ). How can i found out if this converge ?

OpenStudy (anonymous):

Of course it does BECAUSE ...

OpenStudy (anonymous):

There is COMPARISON TEST - YOUR series each summand is smaller than 1/\[n^{2}\] but it is well known that \[\sum_{1}^{\infty} \frac{ 1 }{ n^{2} }\] IS CONVERGENT

OpenStudy (anonymous):

As U realize n* (n+1) \[n(n+1)\]< \[n^{2}\]

OpenStudy (anonymous):

Sorry I meant the sign Bigger n(n+1) > n^2

OpenStudy (anonymous):

So their Inverses are in the opposite Relation !

OpenStudy (anonymous):

Do You understand this argument ?

OpenStudy (anonymous):

n(n+1) is bigger than n^2 so 0 < n^2 < n(n+1). Now with their inverses it reverses it. 0 < 1/[n(n+1)] < 1/[n^2] ? I know that 1/n^2 is convergent so by comparison test I can afirm that 1/ [n(n+1)] is also convergent ? I understood rigth ?

OpenStudy (anonymous):

Yes U DID UNDERSTAND RIGHT. But...

OpenStudy (anonymous):

Writing the answer in such "simplistic" form will show that you do not fully grasp the question. You Should A. READ THE f.-g EXACT THEOREM OF COMAPRISON OF SERIES. B.Understand the importance of "FOR ALL n" in its formulation. C. Stress "for all n" in your written answer

OpenStudy (anonymous):

you can also find the sum if you like

OpenStudy (anonymous):

satellite - this is actually contrary to the best-interest of the student - since he should LEARN the COMPARISON TEST and not some special case which may be USELESS in other examples !!!

OpenStudy (anonymous):

since \(\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}\) your series telescopes

OpenStudy (anonymous):

Satll. No offense intended - just pedagogy